OpenStudy (anonymous):

The ratio of men to women in a certain factory is 3 to 4. There are 222 men. How many workers are there?

OpenStudy (anonymous):

3+4=7 3:7::222:x or \[\frac{ 3 }{ 7 }=\frac{ 222 }{ x },x=\frac{ 222 }{ 3 }\times7=?\]